A particle of charge 3 x 10-9 C is located at point (0,0) in a certain coordinate system. Assuming all lengths are in meters, calculate the potential at the following points: A(3,4) and B(0,2). How much work is needed to take a particle of charge 2 x 10-5 C from point A to point B?








1
Expert's answer
2022-05-02T13:44:12-0400

"U_A=\frac{kq}A=5.4~V,"

"U_B=\frac{kq}B=13.5~V,"

"A=\frac{kqQ}{AB}=150~\mu J."